# An Exploratory Case Study of Query Plan Representations

### v1

@@ A.3 Benchmarking
In this application, we show how to find potential optimization opportunities by comparing the serialized query plans across DBMSs using the unified query plan representation. Query optimization is a critical process for DBMSs’ performance. To evaluate how effective a query optimization is, existing methods depend on measuring DBMSs’ execution time on standard datasets, such as TPC-H (Website, 2023h) and the Join Order Benchmark (JOB) (Leis et al., 2015). The execution time shows the overall performance difference between various DBMSs, but cannot provide possible reasons for performance gaps across DBMSs. UPlan enables comparing various serialized query plans across DBMSs. Specifically, we collected metrics on the number of operations in DBMSs’ query plan representations and show an example of analyzing the difference.
Table 6. The average number of operations in query plans of the queries from the TPC-H benchmark suite.
DBMS Producer Combinator Join Folder Projector Executor Sum
@@
Table 6 shows the average number of operations in each category for the query plans of queries from the TPC-H benchmark suite. We omitted the Consumer category as we did not encounter any such operations. The relational DBMSs, MySQL, PostgreSQL, and TiDB, have more operations than the non-relational DBMSs, MongoDB and Neo4j. It is because relational DBMSs have more operations in the Producer category. To further explain the difference in the Producer category, we looked into query plans and found that each table in relational DBMSs requires at least one operation to read data, while non-relational DBMSs usually read all data in one or two operations. Apart from MongoDB, the other four DBMSs have a similar number of Join operations. The results also show that the query plans of the queries in the TPC-H benchmark suite usually do not cover all categories of operations due to the limited set of queries. The operations of the Producer category are typically expensive for performance, and hence, developers typically aim to analyze and reduce the occurrence of the operations in this category. We explain an example to analyze them as follows.
Figure 4. Variance of the number of Producer operations for each query in TPC-H benchmark across five DBMSs.
Figure 4 shows the variance of the number of operations in the Producer category for each query plan. Among 22 queries, the variances of six queries are more than 5, indicating a significant difference. Queries 2, 8, 5, 7, and 9 have a significant variance due to different data models. For example, for query 2, the DBMSs of relational data models, MySQL, TiDB, and PostgreSQL, have 10, 12, and 9 operations, while the DBMS based on the graph data model, Neo4j, has only 1 operation in the Producer category. Query 11 has a significant variance due to another potential optimization issue, and we explain it as follows.
@@
1SELECT ... FROM PARTSUPP, SUPPLIER, NATION
2WHERE ... GROUP BY ... HAVING ... > (
3 SELECT ... FROM PARTSUPP, SUPPLIER, NATION
4 WHERE ...
5) ORDER BY ...;
6-----------------PostgreSQL-------------------------------------TiDB--------------------
7Combinator->Sort Projector->Project
@@ 28 Producer->Full Table
4 shows query 11 from the TPC-H benchmark suite and the corresponding serialized query plans of PostgreSQL and TiDB using the unified query plan representations in text format. The query references the three tables PARTSUPP, SUPPLIER, NATION twice in the FROM and HAVING clauses respectively. PostgreSQL uses six table scans, one for each table reference in the original query, while TiDB could optimize the query to use only three scans. The table partsupp is scanned twice in lines 21 and 23, because TiDB reduces the data size of table scanning by retrieving a secondary index before the scan operation. The first scan retrieves indexes only to obtain the row id, which is used for the second scan. Suppose both DBMSs execute the operations in order and the same operation has the same performance overhead in both DBMSs, then the query plan with three table scans is more efficient than the query plan with six table scans. UPlan enables this analysis and provides actionable insight for DBMS developers to improve query performance by reducing three repeated table scans. We reported this issue to the PostgreSQL developers, who confirmed that this case indicates an unsupported optimization by PostgreSQL. One developer considered how this could be supported—”[…] maybe there is some easy way to hook this into the same code used by GROUPING SETS […]”.
@@ 6. Discussion
Paths to adoption. Developers can make use of the unified query plan representation by customized converters to convert original serialized query plans into the unified query plan representation. For all three applications we showed, we implemented five customized converters, each of which has around 200 lines of code, within one week only. If the converters can be implemented by DBMS developers or experts, who are well-versed in the query plans, it is plausible that a higher-quality converter could be developed in a shorter time. We hope that DBMSs will directly expose query plans in the unified representation in the long term, thus avoiding a conversion.
@@
Additional use cases. We envision several additional use cases that are enabled by our unified query plan representations. Toward a comprehensive evaluation of query optimization, additional metrics could be explored, such as similarity on tree structures (Yang et al., 2005), to compare different DBMSs’ query plans using our unified query plan representation. To find issues in query optimization, we can apply differential testing (McKeeman, 1998) or other methods to compare the unified query plan representations among DBMSs. Query optimization approaches based on machine learning have been proposed that take query plans as input and output suggestions for indexes (Ding et al., 2019), views (Yuan et al., 2020), and join orders (Marcus and Papaemmanouil, 2018; Yu et al., 2020), so our unified query plan representation would allow exchanging training data in different DBMSs to improve the performance of models.
@@
Threats to validity. Our study faces several threats to validity, which denotes the trustworthiness (Neuman, 1991; Runeson et al., 2012) of the results, and to what extent they are unbiased. A major concern is the degree to which the data and analysis depend on the specific researchers. We followed the best practice of triangulation (Runeson et al., 2012), which refers to taking multiple perspectives toward the same object, to increase reliability. For data triangulation, we collected data from multiple sources: documentation, source code, and third-party applications. For observer triangulation, one author conducted the study, and another author validated the findings against the raw data. For methodological triangulation, we used a qualitative method to analyze query plan representations, and a quantitative method to examine and classify the three conceptual components of query plan representations. Furthermore, we have made the process of data collection and analysis publicly available, along with comprehensive study results presented in the supplementary materials of this paper. Another concern is the degree to which our results can be generalized to and across the query plan representations of other DBMSs. We selected representative DBMSs of various data models: relational, document, graph, and time series. The last concern is the degree to which the study really assesses the research questions we aim at. We collected the data ourselves, and our analysis of the semantics may be inconsistent with the intentions of the developers that implemented the query plans.
7. Related Work
Query plan. The most related work is Substrait (substrait io, 2021), which proposes a well-defined, cross-language specification for data-compute operations on structural data. It was designed for interoperability across different languages and systems. One of its use cases is unifying logical query plan representation for DBMSs. It enables executing the query for a specific DBMS across multiple DBMSs by converting queries into unified logical query plans, which are converted to DBMS-specific physical query plans. However, the unified logical query plans, provided by Substrait, cannot facilitate building general applications based on physical query plans such as demonstrated in this work. Logical query plans represent logical steps to execute a given query, and do not include DBMS-specific internal execution steps, which are required for these applications. In this work, we studied how physical query plans are represented in nine real-world DBMSs and proposed a unified representation for physical query plans.
@@ 7. Related Work
Improving query optimization. Query optimization is a long-standing problem and a core database research topic with a significant body of related work. We discuss several representative works here. Ioannidis et al(Ioannidis, 2003) proposed to sample data in single columns for estimating the cardinality of a query plan. Ilyas et al(Ilyas et al., 2004) proposed to detect correlations across columns, which enables a multi-column sampling (Poosala and Ioannidis, 1997) for a more accurate estimation of cardinality. Lakshmi et al(Lakshmi and Zhou, 1998) proposed using machine learning to estimate cardinality. Estimated cardinality, as well as other estimated costs, are considered in cost models. Babcock et al(Babcock and Chaudhuri, 2005) proposed cost distribution to choose a query plan with the lowest cost in a robust way. Apache published a unified framework Calcite (Begoli et al., 2018) to replace existing query optimizers. Compared to them, we studied various query plan representations, instead of improving query optimization or replacing existing query optimizations.
@@
Benchmark and empirical study of query optimization. Various papers studied query optimizations to understand their performance through empirical studies. Leis et al(Leis et al., 2015) experimentally studied the contributions of query optimization components to overall performance. Ortiz et al(Ortiz et al., 2019) provided an empirical analysis on the accuracy, space, and time trade-off across several machine learning algorithms for cardinality estimation specifically. Harmouch et al(Harmouch and Naumann, 2017) conducted an experimental survey on cardinality estimation, focusing on estimating the number of distinct values. While we also present an empirical study in the context of DBMSs, we focus on query plan representations, rather than the performance, algorithms, and internal components of query optimization.
@@
Applications based on serialized query plans. Several applications based on serialized query plans exist. QE3D (Scheibli et al., 2015) visualizes distributed serialized query plans for an intuitive understanding and analysis. Machine learning algorithms have utilized serialized query plans for query optimization. Yuan et al(Yuan et al., 2020) used machine learning to select optimal views. Yu et al(Yu et al., 2020) used reinforcement learning to determine the join order. Marcus et al(Marcus et al., 2022) and Ryan et al(Marcus et al., 2019) applied machine learning algorithms to generate query plans. Zhao et al(Zhao et al., 2022) used machine learning to convert serialized query plan representations to vector representations to facilitate other machine learning algorithms. In this work, we propose a unified query plan representation, which reduces the effort to build the applications based on serialized query plans.
@@
Standardization. Multiple works were proposed to standardize DBMSs. The most related work is Substrait (substrait io, 2021), which proposes a unified representation for DBMS-agnostic relational algebra, while we studied DBMS-specific query plans. Mitschang (Mitschang, 1988) proposed a unified view of design data and knowledge representation when supporting database systems to non-standard applications, such as Computer-Aided Design (CAD). Ginsburg et al(Ginsburg and Wang, 1992) proposed a unified approach to query sequenced data. Gueidi et al(Gueidi et al., 2021) proposed a unified modeling method for Non-relational DBMSs (NoSQL) to facilitate the applications on NoSQL. Compared to these works, we propose a unified framework for the query plan representations in DBMSs.
8. Conclusion
@@
We have presented an exploratory case study to investigate how query plan representations are in nine widely-used DBMSs. Our study has shown that query plan representations share conceptual components among different DBMSs: operations, properties, and formats. Based on the study, we designed the unified query plan representation to reduce the effort to build applications based on query plans. We implemented a reusable library UPlan to maintain the unified representation, and evaluated it on five DBMSs. The results show that existing testing methods can be efficiently adopted, finding 17 previously unknown and unique bugs. Additionally, existing DBMS-specific visualization tools could support at least five DBMSs by using UPlan with only moderate implementation effort, and UPlan enables comparing query plans in different DBMSs, which provides actionable insights. This paper provides a comprehensive study of query plan representations, and can be used as a reference for other research on serialized query plans. We believe the unified query plan representation provides more opportunities to research serialized query plans in the future.

### v3

@@ A.3 Benchmarking
In this application, we show how to find potential optimization opportunities by comparing the serialized query plans across DBMSs using the unified query plan representation. Query optimization is a critical process for DBMSs’ performance. To evaluate how effective a query optimization is, existing methods depend on measuring DBMSs’ execution time on standard datasets, such as TPC-H [45] and the Join Order Benchmark (JOB) [52]. The execution time shows the overall performance difference between various DBMSs, but cannot provide possible reasons for performance gaps across DBMSs. UPlan enables comparing various serialized query plans across DBMSs. Specifically, we collected metrics on the number of operations in DBMSs’ query plan representations and show an example of analyzing the difference.
TABLE VI: Average number of operations in query plans from TPC-H.
DBMS Prod. Comb. Join Folder Proj. Exec. Sum
@@
TABLE VII: Average number of operations in query plans from YCSB for MongoDB and WDBench for Neo4j.
DBMS Prod. Comb. Join Folder Proj. Exec. Sum
MongoDB 1.00 0.00 0.00 0.00 0.00 0.00 1.00
Neo4j 0.45 0.00 3.73 0.00 0.37 1.44 5.99
Table VI shows the average number of operations in each category for the query plans of queries from the TPC-H benchmark. We omitted the Consumer category as we did not encounter any such operations. Overall, the relational DBMSs, MySQL, PostgreSQL, and TiDB, have more operations than the non-relational DBMSs, MongoDB and Neo4j. It is because relational DBMSs have more operations in the Producer category. To further explain the difference in the Producer category, we looked into query plans and found that each table in relational DBMSs requires at least one operation to read data, while non-relational DBMSs usually read all data in one or two operations. Apart from MongoDB, the other four DBMSs have a similar number of Join operations. The results also show that the query plans of the queries in the TPC-H benchmark suite usually do not cover all categories of operations due to the limited set of queries.
Similarly, Table VII shows the average number of operations for the query plans from the YCSB and WDBench benchmarks respectively. Both benchmarks are designed specifically for NoSQL and graph DBMSs, but mainly consider input diversity instead of internal execution diversity, as query plans are DBMS-specific and challenging to examine. As a result, compared to TPC-H, YCSB and WDBench have a similar distribution of operations. For some categories, the query plans in YCSB and WDBench include even fewer operations than those in TPC-H. For example, YCSB does not expose any operation in the Projection category and WDBench does not expose any operation in the Combinator and Folder categories. This shows that the TPC-H benchmark can efficiently expose operations in query plans.
Figure 4: Variance of the number of Producer operations for each query in TPC-H benchmark across five DBMSs.
To find potential optimization opportunities, we analyzed the operations of the Producer. Across the five DBMSs, the operations of the Producer category are typically expensive for performance, and hence, developers typically aim to analyze and reduce the occurrence of the operations in this category. Figure 4 shows the variance of the number of operations in the Producer category for each query plan. Among 22 queries, the variances of six queries are more than 5, indicating a significant difference. Queries 2, 8, 5, 7, and 9 have a significant variance due to different data models. For example, for query 2, the DBMSs of relational data models, MySQL, TiDB, and PostgreSQL, have 10, 12, and 9 operations, while the DBMS based on the graph data model, Neo4j, has only 1 operation in the Producer category. Query 11 has a significant variance due to a potential optimization issue, and we explain it as follows.
@@
1SELECT ... FROM PARTSUPP, SUPPLIER, NATION WHERE ...
2HAVING ... > (SELECT ... FROM PARTSUPP, SUPPLIER, NATION WHERE ...) ...;
3--------------------------------------------------------
4PostgreSQL: TiDB:
5Combinator->Sort Projector->Project
@@ 26 Producer->Full Table
4 shows query 11 from the TPC-H benchmark suite and the corresponding serialized query plans of PostgreSQL and TiDB using the unified query plan representations in text format. The query references the three tables PARTSUPP, SUPPLIER, NATION twice in the FROM and HAVING clauses respectively. PostgreSQL uses six table scans, one for each table reference in the original query, while TiDB could optimize the query to use only three scans. The table partsupp is scanned twice in lines 19 and 21, because TiDB reduces the data size of table scanning by retrieving a secondary index before the scan operation. The first scan retrieves indexes only to obtain the row id, which is used for the second scan. Suppose both DBMSs execute the operations in order and the same operation has the same performance overhead in both DBMSs, then the query plan with three table scans is more efficient than the query plan with six table scans.
We quantitatively evaluated the potential performance improvement contributed by our analysis. It is challenging to modify PostgreSQL source code to more effectively optimize this query plan. To estimate the performance improvement, we evaluated the actual execution time for the six table scans in PostgreSQL’s query plan using the SQL prefix EXPLAIN ANALYZE, which executes the query and returns the execution information of the query plan including the execution time for each operation. For 10 GB data of TPC-H benchmark, the whole query consumes 1503ms, and the six table scans consume 214ms, 5ms, 1ms, 216ms, 2ms, and 189ms, respectively. If PostgreSQL optimizes this query plan to remove the last three table scans, the execution time would be significantly reduced by 407ms ($216+2+189$), which accounts for 27% of the overall execution time of the query. Note that we assumed that eliminating the three operations would not influence the execution time of other operators. UPlan enables this analysis and provides actionable insight for DBMS developers to improve query performance by reducing three repeated table scans. We reported this issue to the PostgreSQL developers, who confirmed that this case indicates an unsupported optimization by PostgreSQL. One developer considered how this could be supported—”[…] maybe there is some easy way to hook this into the same code used by GROUPING SETS […]”.
@@ VI Discussion
Paths to adoption. Developers can use the unified query plan representation by customized converters to convert original serialized query plans into the unified query plan representation. For all three applications we showed, we implemented five customized converters, each of which has around 200 lines of code, within one week only. If the converters can be implemented by DBMS developers or experts, who are well-versed in the query plans, it is plausible that a higher-quality converter could be developed in a shorter time. We hope that DBMSs will directly expose query plans in the unified representation in the long term, thus avoiding a conversion.
@@
Additional use cases. We envision several additional use cases that are enabled by our unified query plan representations. Toward a comprehensive evaluation of query optimization, additional metrics could be explored, such as similarity on tree structures [53], to compare different DBMSs’ query plans using our unified query plan representation. To find issues in query optimization, we can apply differential testing [54] or other methods to compare the unified query plan representations among DBMSs. Query optimization approaches based on machine learning have been proposed that take query plans as input and output suggestions for indexes [55], views [56], and join orders [57, 58], so our unified query plan representation would allow exchanging training data in different DBMSs to improve the performance of models.
Substrait. UPlan complements Substrait [59], which provides a DBMS-agnostic specification for logical query plan representation to enable cross-language serialization for relational algebra. First, Substrait mainly focuses on logical query plans, while UPlan focuses on physical query plans. Many physical operations are DBMS-specific and are not considered by Substrait. In Section III, our study based on real-world query plans shows that the physical operations of Executor and Consumer do not have corresponding operators in relational algebra, and Substrait does not support them. Integrating UPlan as a backend component for Substrait would allow it to analyze, optimize, and test DBMS-specific query executions comprehensively. Second, UPlan enables different application domains compared to Substrait. UPlan aims to reduce the effort of building applications on query plans instead of cross-language serialization for relational algebra. The unified logical query plans, provided by Substrait, cannot facilitate building general applications based on physical query plans such as demonstrated in this work. Lastly, UPlan offers easier integration with real-world DBMSs. Unlike Substrait, which requires additional support for DBMSs to both output logical query plans in Substrait format and execute those plans, UPlan directly parses existing query plans without requiring any modifications to the DBMS. This significantly reduces the implementation effort typically needed by Substrait.
Completeness. We believe our study and implementation are sufficiently complete as we studied query plans from various sources. As we observed in our study, DBMSs lack detailed and standard documents for query plans, so we comprehensively studied query plans from documents, source code, and concrete executions with the workloads from TPC-H, WDBench, and YCSB benchmarks. We also evaluated our implementation of UPlan on visualization and benchmarking with three benchmarks. However, the possibility of missing features in query plans still exists. For example, some properties may be dynamically generated without detailed documentation. To observe them, we require specific workloads, which we may overlooked. To alleviate this potential issue, we designed an extensible UPlan. Such missing properties can be easily integrated into the unified query plan representation by adding keyword definitions as we explained in Section IV.
@@
Threats to validity. Our study faces several threats to validity, which denotes the trustworthiness [60, 12] of the results, and to what extent they are unbiased. A major concern is the degree to which the data and analysis depend on the specific researchers. We followed the best practice of triangulation [12], which refers to taking multiple perspectives toward the same object, to increase reliability. For data triangulation, we collected data from multiple sources: documentation, source code, and third-party applications. For observer triangulation, one author conducted the study, and another author validated the findings against the raw data. For methodological triangulation, we used a qualitative method to analyze query plan representations, and a quantitative method to examine and classify the three conceptual components of query plan representations. Furthermore, we have made the process of data collection and analysis publicly available, along with comprehensive study results presented in the supplementary materials of this paper. Another concern is the degree to which our results can be generalized to and across the query plan representations of other DBMSs. We selected representative DBMSs of various data models: relational, document, graph, and time series. The last concern is the degree to which the study really assesses the research questions we aim at. We collected the data ourselves, and our analysis of the semantics may be inconsistent with the intentions of the developers that implemented the query plans.
VII Related Work
@@
Improving query optimization. Query optimization is a long-standing challenge and a core database research topic with a significant body of related work. We discuss several representative works here. Ioannidis et al[32] proposed to sample data in single columns for estimating the cardinality of a query plan. Ilyas et al[61] proposed to detect correlations across columns, which enables a multi-column sampling [62] for a more accurate estimation of cardinality. Lakshmi et al[63] proposed using machine learning to estimate cardinality. Estimated cardinality, as well as other estimated costs, are considered in cost models. Babcock et al[64] proposed cost distribution to choose a query plan with the lowest cost in a robust way. Apache published a unified framework Calcite [65] to replace existing query optimizers. Compared to them, we studied various query plan representations, rather than improving query optimization.
@@
Benchmark and empirical study of query optimization. Various papers studied query optimizations to understand their performance through empirical studies. Leis et al[52] experimentally studied the contributions of query optimization components to overall performance. Ortiz et al[66] provided an empirical analysis on the accuracy, space, and time trade-off across several machine learning algorithms for cardinality estimation specifically. Harmouch et al[67] conducted an experimental survey on cardinality estimation, focusing on estimating the number of distinct values. While we also present an empirical study in the context of DBMSs, we focus on query plan representations, rather than the performance, algorithms, and internal components of query optimization.
@@
Applications based on serialized query plans. Several applications based on serialized query plans exist. QE3D [68] visualizes distributed serialized query plans for an intuitive understanding and analysis. Machine learning algorithms have utilized serialized query plans for query optimization. Yuan et al[56] used machine learning to select optimal views. Yu et al[58] used reinforcement learning to determine the join order. Marcus et al[69] and Ryan et al[70] applied machine learning algorithms to generate query plans. Zhao et al[71] used machine learning to convert serialized query plan representations to vector representations to facilitate other machine learning algorithms. In this work, we propose a unified query plan representation, which reduces the effort to build the applications based on serialized query plans.
@@
Standardization. Multiple works were proposed to standardize DBMSs. The most related work is Substrait [59], which proposes a unified representation for DBMS-agnostic relational algebra, while we studied DBMS-specific query plans. Feng et al[72] proposed a unified architecture to reduce the implementation effort for in-RDBMS analytics. Mitschang [73] proposed a unified view of design data and knowledge representation when supporting database systems to non-standard applications, such as Computer-Aided Design (CAD). Ginsburg et al[74] proposed a unified approach to query sequenced data. Gueidi et al[75] proposed a unified modeling method for Non-relational DBMSs (NoSQL) to facilitate the applications on NoSQL. Compared to these works, we propose a unified framework for the query plan representations in DBMSs.
VIII Conclusion
@@
We have presented an exploratory case study to investigate how query plan representations are in nine widely-used DBMSs. Our study has shown that query plan representations share conceptual components among different DBMSs: operations, properties, and formats. Based on the study, we designed the unified query plan representation to reduce the effort to build applications based on query plans. We implemented a reusable library UPlan to maintain the unified representation, and evaluated it on five DBMSs. The results show that existing testing methods can be efficiently adopted, finding 17 previously unknown and unique bugs. Additionally, existing DBMS-specific visualization tools could support at least five DBMSs by using UPlan with only moderate implementation effort, and UPlan enables comparing query plans in different DBMSs, which provides actionable insights. This paper provides a comprehensive study of query plan representations, and can be used as a reference for other research on query plans. We believe the unified query plan representation provides more opportunities to research serialized query plans in the future.
Acknowledgments
This research is supported by the National Research Foundation, Singapore, and Cyber Security Agency of Singapore under its National Cybersecurity R&D Programme (Fuzz Testing). Any opinions, findings and conclusions, or recommendations expressed in this material are those of the author(s) and do not reflect the views of National Research Foundation, Singapore, and Cyber Security Agency of Singapore.